# Learning Unmasking Policies for Diffusion Language Models

### v2

@@ A.1 Figure 1 replicated for {LLaDA-8B-Instruct, Dream-7b-Instruct} $\times$ {GSM8k, MATH-500}
A.2 Figure 3 replicated for Dream-7b-Instruct
@@
Figure 7: Results for Dream in semi-AR (Figure˜7(a) & Figure˜7(c)) and full-diffusion (Figure˜7(b) & Figure˜7(d)) generation regimes. For the policies we vary $\alpha\in\{10,3,1,0.3,0\}$ and use $\tau_{\pi}=0.5$ for $BL=32$ and $\tau_{\pi}=1$ for $BL=256$. Unlike for LLaDA, where the learned policies clearly match Fast-dLLM in the mid-to-high NFEs regime for semi-AR generation, a small gap appears to remain for Dream. Furthermore, we found that the $\alpha=10$ policy did not converge in the $BL=32$ setting when using Dream as the underlying language model (downstream accuracy thus not shown in (a), (c), to reduce clutter in the figures). Meanwhile, in the full-diffusion setting we do observe some performance gains, but here too they are smaller than were seen for LLaDA.
A.3 Finer grid for $\alpha$
@@
Figure 8: $BL=32$ results for LLaDA with a denser regularization grid, $\alpha\in\{10.0,9.0,\ldots,1.0,0.3,0.0\}$. Single training seed due to cost; error bars show (min, max) over three test-time seeds. Note that for $\alpha\geq 4.0$, the change in NFEs is not monotonic; different values lead to convergence either close to the $\alpha=3.0$ policy, or to that of $\alpha=10.0$.
A.4 Impact of policy temperature $\tau_{\pi}$
@@
Figure 9: We study the effect of changing the policy temperature $\tau_{\pi}$ (cf. Section˜3.2). For each $\alpha\in{3,0.3,0}$, we construct a corresponding test-time Pareto frontier by varying $\tau_{\pi}\in{1.5,1.0,0.5}$. Interestingly, in some cases—such as $\alpha=0$ with BL=32BL=32—adjusting $\tau_{\pi}$ enables an effective trade-off between compute and performance. Moreover, we find that $\tau_{\pi}=0.5$ is optimal in the semi-AR, while $\tau_{\pi}=1$ performs best in the non-semi-AR ($BL=256$) setting.
A.5 Model transfer results
@@
Figure 10: Model transfer results. We use policies trained on LLaDA and evaluate them on Dream with $\tau_{\pi}=0.5$. Encouragingly, transferred policies give almost identical results compared to training Dream specific policies (cf. Figure˜7(a) & Figure˜7(c)). Note that the $\alpha=10$ policy is represented separately by a ( ) in the lower left to avoid misleading visualization when interpolating to $\alpha=3$.
A.6 Sequence-length transferability
@@
Figure 11: $BL=L=256$-trained policies from Section˜4.2 evaluated with a 2x longer sequence length ($BL=L=512$) with $\tau_{\pi}=1$. Note that the learned policies yield almost identical performance, while the heuristic methods degrade further compared to $L=256$ (cf. Figure˜3(b) & Figure˜3(d)). Both for LLaDA-8B-Instruct.
Appendix B Training and Policy Network Configuration
Table 1: Training and policy configuration for our main experiments.
Category Parameter Value
Training Learning rate 3e-5
LR scheduler Cosine
Warmup steps 100
Effective batch size 16
Weight decay 0.1
Max gradient norm 0.2
Clipping factor ($\epsilon$) 0.5 (0.2 for expert steering)
GRPO Training data GSM8k and MATH training set mixture
Num train epochs 1
Group size 8
$\beta$ (KL penalty) 0.0
Policy Network Number of transformer blocks 1
Hidden dimension 128
Feedforward dimension 512
Number of heads 2
Time embedding dim 128
Total parameter count 300K
Appendix C Dynamic Plackett-Luce Sampling
@@
Appendix D Expert Steering
As mentioned in Section˜4.2, we find that naïvely training confidence policies directly on the full-length generation ($BL=L$) task yields policies which beat out the heuristic methods, but underperform the ones trained in the semi-AR setting (cf. Figure˜3).
Since semi-AR decoding lies within the function class representable by the policy $\pi_{\phi}$—that is, there is some $\phi$ such that $\pi_{\phi}$ approximates the semi-AR setting—we hypothesize that our failure to obtain similar policies is due to the vanishingly small probability of encountering autoregressive-like rollouts by chance. More concretely, image trying to train a policy on a task for which purely AR generation is substantially more effective than any other decoding strategy. In such a scenario, starting from a randomly initialized policy, the probability of observing a rollout resembling a purely AR sequence is approximately $1/L!\approx 0$ (where $L$ is the generation length), making it extremely unlikely to sample such rollouts during RL training.
@@ Appendix D Expert Steering
Appendix E Extended Background
E.1 Discrete diffusion variants and continuous relaxations
@@
with $\alpha_{t}\in[0,1]$, ${\bm{e}}_{m}$ denotes a one-hot vector for $\bm{M}$} token, and $\mathbf{1}$ represents a vector of all 1s of size $V$, and (ii) uniform-state diffusion, which replaces tokens uniformly at random,
$Q_{t}^{\text{uni}}\;=\;(1-\alpha_{t})I\;+\;\alpha\,\tfrac{\mathbf{1}\mathbf{1}^{\top}}{V}.$
@@
A complementary approach to discrete-state diffusion is to keep the diffusion process continuous in both time and input by operating on token embeddings. For example in DBLP:journals/corr/abs-2211-15089, an orthogonal line embeds tokens into $\mathbb{R}^{d}$ and applies fully continuous diffusion in both time and input space. Training can be done with cross-entropy via score interpolation, and sampling uses ODE/SDE solvers with tools like classifier-free guidance. This preserves the continuous machinery but introduces an embedding decoding interface and relaxes exact discreteness during the trajectory.

### v3

@@ B.1 Figure 1 replicated for {LLaDA-8B-Instruct, Dream-7b-Instruct} $\times$ {GSM8k, MATH-500}
B.2 Figure 4 replicated using wall-clock time as the efficiency measure
(a) LLaDA GSM8k
(b) LLaDA GSM8k
(c) LLaDA MATH
(d) LLaDA MATH
Figure 10: Figure˜4 reproduced using wall-clock time (in seconds) as the efficiency metric instead of NFEs. The difference in our policy curves ( ) when measured in wall-clock time versus NFEs (see Figure˜4) is minimal to non-existent, demonstrating the negligible computational overhead of our policy. This low overhead is primarily due to the small size of the unmasking model relative to the base dLLM (300K vs. 8B parameters in the case of LLaDA). All experiments run on A100 GPUs.
B.3 Figure 4 replicated for Dream-7b-Instruct
@@
Figure 11: Results for Dream in semi-AR (Figure˜11(a) & Figure˜11(c)) and full-diffusion (Figure˜11(b) & Figure˜11(d)) generation regimes. For the policies we vary $\alpha\in\{10,3,1,0.3,0\}$ and use $\tau_{\pi}=0.5$ for $BL=32$ and $\tau_{\pi}=1$ for $BL=256$.
B.4 Controllability of learning unmasking policies with RL (via $\alpha$)
@@
Figure 12: $BL=32$ results for LLaDA with a denser regularization grid, $\alpha\in\{10.0,9.0,\ldots,1.0,0.3,0.0\}$. Single training seed due to cost; error bars show (min, max) over three test-time seeds. Note that for $\alpha\geq 4.0$, the change in NFEs is not monotonic; different values lead to convergence either close to the $\alpha=3.0$ policy, or to that of $\alpha=10.0$. Points for policy results are connected by $\alpha$ to emphasize the non-monotone behavior.
(a) GSM8k, $BL=256$
(b) MATH-500, $BL=32$
Figure 13: $BL=256$ (GSM8k) and $BL=32$ (MATH-500) results for LLaDA with all training seeds scattered. Even for a fixed value of $\alpha$, the resulting policy can vary in accuracy and speed due to the randomness of the training procedure. Marker shape denotes the value of $\alpha$.
B.5 Impact of policy temperature $\tau_{\pi}$
@@
Figure 14: We study the effect of changing the policy temperature $\tau_{\pi}$ (cf. Section˜3.2). For each $\alpha\in\{3,0.3,0\}$, we construct a corresponding test-time Pareto frontier by varying $\tau_{\pi}\in\{1.5,1.0,0.5\}$. Interestingly, in some cases—such as $\alpha=0$ with BL=32BL=32—adjusting $\tau_{\pi}$ enables an effective trade-off between compute and performance. Moreover, we find that $\tau_{\pi}=0.5$ is optimal in the semi-AR, while $\tau_{\pi}=1$ performs best in the full-diffusion ($BL=L=256$) setting.
B.6 Model transfer results
@@
Figure 15: Model transfer results. We use policies trained on LLaDA and evaluate them on Dream with $\tau_{\pi}=0.5$. Encouragingly, transferred policies give comparable results to training Dream specific policies (cf. Figure˜11(a) & Figure˜11(c)). Note that the $\alpha=10$ policy is represented separately by a ( ) in the lower left to avoid misleading visualization when interpolating to $\alpha=3$.
B.7 Sequence-length transferability
@@
Figure 16: $BL=L=256$-trained policies from Section˜4.2 evaluated with a 2x longer sequence length ($BL=L=512$) with $\tau_{\pi}=1$. Note that the learned policies yield almost identical performance, while the heuristic methods degrade further compared to $L=256$ (cf. Figure˜4(b) & Figure˜4(d)). Both for LLaDA-8B-Instruct.
Appendix C What are RL unmasking policies actually doing?
(a) $BL=32$
(b) $BL=256$
Figure 17: Results for LLaDA in semi-AR ($BL=32$; replicated from Figure˜4(a)) and full-diffusion ($BL=256$; figure replicated from Figure˜4(b)) generation regimes. We highlight the policies used in our qualitative analysis of unmasking orders:   for BL32-slow,   for BL32-fast, and   for BL256.
To better understand the behaviour of our learned policies, we conduct a qualitative analysis comparing generation/unmasking trajectories produced by our RL policies against a confidence heuristic like Fast-dLLM (fastdllm2025). We consider three settings that span different points on the accuracy-efficiency Pareto frontier and different generation regimes (see Figure˜17):
  • BL32-slow ( ): Semi-AR ($BL=32$) with Fast-dLLM ($\lambda=0.9$) vs. Policy ($\alpha=0.3$)
  • BL32-fast ( ): Semi-AR ($BL=32$) with Fast-dLLM ($\lambda=0.1$) vs. Policy ($\alpha=10$)
  • BL256 ( ): Full-diffusion ($BL=L=256$) with Fast-dLLM ($\lambda=0.9$) vs. Policy ($\alpha=0.3$) vs. Policy with expert steering ($\alpha=0.3$)
Given a prompt $\bm{x}$} and generated answer $\hat{{\bm{y}}}$ (with a maximum answer length and maximum number of steps to set to $L$ and $T$, respectively), the unmasking order [^3] Note that we consider partial orders here, as multiple tokens can be unmasked simultaneously. is defined as ${\bm{o}}\in[T]^{L}$, where $o_{k}:=\min\{t\mid y_{t}^{k}\neq{\bm{M}}\}$ denotes the timestep when the token at position $k$ is first unmasked. For presentation clarity, we reverse the time axis relative to the main paper, so time flows as $t=1,2,\ldots,T$. For all experiments, we use $N=100$ GSM8k test samples with $L=T=256$.
We describe main insights for each setting next.
C.1 BL32-slow
Figure 18: Spearman rank correlation between the unmasking orders of Fast-dLLM and our RL policies per block for semi-AR policies considered here. Note that for certain blocks, the unmasking orders exhibit zero correlation: first and last block for BL32-fast( ), and last block for BL32-slow ( ).
In this setting, our policy achieves performance nearly identical to Fast-dLLM in both accuracy and efficiency (see   in Figure˜17(a)). Despite these similarities, we find that the learned policy adopts a somewhat different sampling strategy compared to plain confidence thresholding. This is visually apparent in Figure 19, where we observe that, although both strategies predominantly follow a left-to-right generation pattern (due to semi-AR regime), the policy’s unmasking order lines ( ) display more "wiggliness" within each block. This is due to the fact that, as shown in Figure˜20, the policy is far less likely to unmask adjacent tokens simultaneously—doing so for only around $\sim 20\%$ of tokens, in contrast to Fast-dLLM, which does so for $\sim 65\%$ on average. This suggests that the policy learns a more dispersed unmasking strategy, favoring token commitments at scattered positions within each block rather than in contiguous chunks. Such a scattered approach may enable the model to gather more diverse contextual information before finalizing neighboring tokens, potentially mitigating errors that arise from parallel generation. The more scattered generation pattern within each block is also evident in the token generation trajectories shown in Figure˜23.
Figure 19: BL32-slow: unmasking orders for policy sampling (${\bm{o}}_{\alpha=0.3}$) and Fast-dLLM (${\bm{o}}_{\lambda=0.9}$) for four GSM8k test samples under semi-AR generation ($BL=32,L=256$). Vertical dashed lines indicate the position where the model outputs the end-of-sequence (EOS) token. Interestingly, Fast-dLLM appears more effective at accelerating generation after producing the EOS token, whereas policy sampling tends to sometimes waste compute by not unmasking all padding tokens (within the current block) after EOS in parallel, see the third test sample.
The two methods also differ substantially in how they allocate compute across blocks. As shown in Figure˜5(a), Fast-dLLM assigns a larger number of NFEs to earlier blocks, allocating $\sim\!17$ NFEs to the first block and gradually decreasing to just $\sim\!4$ NFEs in the final block. In contrast, our learned policy distributes compute almost uniformly across blocks, using $\sim\!10$ NFEs per block. This likely also contributes to the decreasing Spearman rank correlation of unmasking orders between the policy and Fast-dLLM sampling across blocks (see Figure˜18).
Figure 20: BL32-slow: for each of the $N=100$ samples, we report the frequency of tokens whose right-adjacent token was unmasked at the same sampling step under Fast-dLLM versus our policy sampling. Interestingly, policy sampling unmasks adjacent tokens much less frequently than Fast-dLLM.
C.2 BL32-fast
Here both strategies achieve very fast but low-accuracy sampling (see   in Figure˜17(a)). However, policy sampling significantly outperforms Fast-dLLM, achieving approximately $\sim 38\%$ accuracy compared to $\sim 18\%$. Examining the unmasking orders in Figure˜21 and the average NFEs per block in Figure˜5(b), we observe that policy sampling tends to slow down in the final block, typically generating the numerical answer last (see also the token generation trajectory in Figure˜25 and how the tokens of the numerical answer are generated last under policy sampling). This behavior may help explain its improved performance over Fast-dLLM and shows the promise of RL for discovering new sampling techniques.
C.3 BL256
Lastly, we consider the full-diffusion setting ($BL=L=256$). While all three strategies achieve a comparable number of NFEs (see   in Figure˜17(b)), their accuracy differs significantly. The fact that these strategies implement notably different sampling procedures is further evidenced by the low values of Spearman rank correlations $0.26\pm 0.21$ when comparing Fast-dLLM orders with policy sampling, and $0.24\pm 0.51$ when comparing Fast-dLLM with policy trained with expert steering. These differences are also reflected in the divergent unmasking orders shown in Figure˜22 (also refer to Figure˜6 for a plot of average unmasking order across samples).
Figure 21: BL32-fast: unmasking orders for policy sampling (${\bm{o}}_{\alpha=10}$) and Fast-dLLM (${\bm{o}}_{\lambda=0.1}$) for four GSM8k test samples under semi-AR generation ($BL=32,L=256$). Under low-NFE constraints, both strategies often predict all tokens within a block simultaneously (as indicated by flat unmasking-order lines). In the final block, when generating the numerical answer, the policy slows down.
Starting with Fast-dLLM ( ), we observe in Figure˜22 that it often generates tokens in a "reverse" order (right-to-left), see also the token trajectories in Figure˜28). This behavior likely stems from the fact that LLaDA did not exclude padding tokens from the loss computation during SFT (llada2025) to preserve the dLLM’s ability to generate sequences of varying (effective) lengths. Consequently, the model overfits to padding tokens, which confidence-based methods like Fast-dLLM then mistakenly prioritize, leading to their early generation. This helps explain the poor performance of confidence-based strategies observed in Figure˜1, highlighting the fragility of heuristic-driven approaches.
Looking at the unmasking orders of policy sampling ( ), the policy appears to implement a near-random unmasking strategy, as there is little evident spatial structure in the unmasking order. This is encouraging, as it suggests that the policy can learn to disregard corrupted signals—such as inflated confidence for padding tokens—without explicit supervision, leading to more robust performance across different generation settings (semi-AR vs full-diffusion). Lastly, we observe that the policy with expert steering ( ) generates in an AR fashion (left-to-right). This behavior likely accounts for its superior performance in the full-diffusion setting. It indicates that, during full-diffusion training, the policy is able to pick up the accuracy benefits of AR generation from its expert demonstrations—drawn from the semi-AR regime (Appendix F)—while still learning to generate significantly faster than a purely AR strategy that unmaskes one token at a time.
Figure 22: BL256: unmasking orders for policy sampling (${\bm{o}}_{\alpha}=0.3$) and Fast-dLLM (${\bm{o}}_{\lambda=0.9}$) for four GSM8k test samples under full-diffusion generation ($BL=L=256$).
Appendix D Generation Trajectories Visualizations
Here we visualize token generation trajectories for the various sampling strategies considered in our work. Specifically, for a randomly selected GSM8k test sample, we display the generated token at each of the $L=256$ positions, along with its unmasking time (shown in the bottom right corner of each cell, starting from $t=0$). For visual clarity, token cells are color-coded according to their unmasking time, with blue indicating earlier and red indicating later unmasking. Note that for semi-AR generations, color-coding is done within each block separately.
Figure 23: BL32-slow: generation trajectory for policy sampling ($\alpha=0.3$, semi-AR).
Figure 24: BL32-slow: generation trajectory for Fast-dLLM sampling ($\lambda=0.9$, semi-AR).
Figure 25: BL32-fast: generation trajectory for policy sampling ($\alpha=10$, semi-AR).
Figure 26: BL32-fast: generation trajectory for Fast-dLLM sampling ($\lambda=0.1$, semi-AR).
Figure 27: BL256: generation trajectory for policy sampling ($\alpha=0.3$, full-diffusion).
Figure 28: BL256: generation trajectory for Fast-dLLM sampling ($\lambda=0.9$, full-diffusion).
Figure 29: BL256: generation trajectory for policy sampling ($\alpha=0.3$, expert steering, full-diffusion).
Appendix E Dynamic Plackett-Luce Sampling
@@
Appendix F Expert Steering
As mentioned in Section˜4.2, we find that naïvely training confidence policies directly on full-diffusion generation ($BL=L$) task yields policies which beat out the heuristic methods, but underperform the ones trained in the semi-AR setting (cf. Figure˜4).
Since semi-AR decoding lies within the function class representable by the policy $\pi_{\phi}$—that is, there is some $\phi$ such that $\pi_{\phi}$ approximates the semi-AR setting—we hypothesize that our failure to obtain similar policies is due to the vanishingly small probability of encountering autoregressive-like rollouts by chance. More concretely, imagine trying to train a policy on a task for which purely AR generation is substantially more effective than any other decoding strategy. In such a scenario, starting from a randomly initialized policy, the probability of observing a rollout resembling a purely AR sequence is approximately $1/L!\approx 0$ (where $L$ is the generation length), making it extremely unlikely to sample such rollouts during RL training.
@@ Appendix F Expert Steering
Appendix G Extended Background
@@
with $\alpha_{t}\in[0,1]$, ${\bm{e}}_{m}$ denotes a one-hot vector (with dimension $V$) for $\bm{M}$} token, and $\mathbf{1}$ represents a vector of all 1s of size $V$, and (ii) uniform-state diffusion, which replaces tokens uniformly at random,
$Q_{t}^{\text{uni}}\;=\;(1-\alpha_{t})I\;+\;\frac{\alpha_{t}}{V}\,\mathbf{1}\mathbf{1}^{\top}.$
@@
A complementary approach to discrete-state diffusion is to keep the diffusion process continuous in both time and input by operating on token embeddings. For example in dieleman2022continuous, an orthogonal line embeds tokens into $\mathbb{R}^{d}$ and applies fully continuous diffusion in both time and input space. Training can be done with cross-entropy via score interpolation, and sampling uses ODE/SDE solvers with tools like classifier-free guidance. This preserves the continuous machinery but introduces an embedding decoding interface and relaxes exact discreteness during the trajectory.
@@
Appendix H Training and Policy Network Configuration
Table 1: Training and policy configuration for our main experiments.
Category Parameter Value
Training Learning rate 3e-5
LR scheduler Cosine
Warmup steps 100
Effective batch size 16
Weight decay 0.1
Max gradient norm 0.2
Clipping factor ($\epsilon$) 0.5 (0.2 for expert steering)
GRPO Training data GSM8k and MATH training set mixture
Num train epochs 1
Group size 8
$\beta$ (KL penalty) 0.0
Correctness reward $r({\bm{y}},{\bm{y}}_{t})$ We use d1 (zhao2025d1) rewards.
Policy Network Number of transformer blocks 1
Hidden dimension 128
Feedforward dimension 512
Number of heads 2
Time embedding dim 128
Total parameter count 300K
Appendix I Policy Architecture Diagram
[…]